# Preprint version

### v2

@@ 1 Introduction
Deep learning models have enabled significant improvements in Time Series (TS) classification, due to their powerful representational capabilities [1]. However, most models suffer from a lack of generalization and can hardly be transferred from their training domain (the so-called source domain) to an unfamiliar, external domain (the target domain) [2]. This limitation is mainly caused by distribution shifts, a very common phenomenon in TS [3, 4]. To address this problem, Unsupervised Domain Adaptation (UDA) attempts to find a domain-invariant feature space for labeled source samples and unlabeled target samples. Recently, several UDA algorithms have been benchmarked for TS tasks [2], mainly using image processing architectures. The limited performance of UDA methods on TS datasets compared to their high accuracy on image tasks underscores the need for TS-specific architectures. Since [2], many approaches have enhanced the performance of TS representation by identifying invariant temporal features specifically tailored for TS [5, 6, 7, 8, 9]. For example, frequency-based features extracted from neural operators, such as Fourier Neural Operators (FNO) [10], have improved domain-invariant representations and have been used for UDA on TS [11].
@@
The UDA framework actually covers a special case where the classes are the same in both domains (despite some small drift between domains) known as the closed-set assumption. Motivated by this limitation, Universal Domain Adaptation (UniDA) was introduced [12]. In UniDA, the target domain remains unlabeled, certain classes are shared between the source and target domains, while others are exclusive to either domain. Unlike standard UDA, UniDA helps to discover unknown class samples that are unique to the target domain. This capability is crucial in TS classification, where identifying unseen patterns is often critical [13].
@@
UniDA encompasses two primary tasks: alignment of common classes across domains and isolation of unknown Out-Of-Distribution (OOD) samples. Several approaches have been developed to address both tasks using different strategies. The alignment module is usually a pre-existing UDA method that is slightly modified to account for unknown labels. For instance, UAN [12] uses an adversarial UDA approach introduced in [14, 15], while DANCE [16] adopts a similarity-based clustering approach inspired by [17]. Similarly, PPOT [18] and UniOT [19] align source and target common samples using Optimal Transport (OT) [20], as proposed in [21, 22]. Alternatively, UniAM [23] introduces a novel alignment framework leveraging sparse data representation and Vision Transformer (ViT) architectures.
@@
On the other hand, the cited approaches rely on OOD detection methods that are highly threshold-dependent. These methods aim to separate common from unknown samples by thresholding discriminability metrics such as entropy [16, 12], OT masses [18, 19], reconstruction error [23], or combinations of metrics [24]. The threshold value is typically either a fine-tuned hyperparameter [12, 16, 18, 23] or a fixed value [19] determined from a dataset. However, as shown in experiments in Section 4, the optimal threshold often varies between datasets and across tasks. Nevertheless, most previous approaches rely on robust state-of-the-art alignment modules to address the inefficiencies in OOD detection caused by rigid threshold definitions. These approaches fail to consider that inadequate OOD detection can compromise the performance of the alignment module. In consequence, these observations highlight the need for a dynamic threshold selection to better adapt UniDA methods to different datasets.
@@
Few approaches attempt to address the dynamic nature of the threshold value. For example, TNT [25] computes a task-specific threshold, but this computation is based solely on the source dataset. Domain shifts may render such computations ineffective on the target dataset. Moreover, OVANet [26] employs a One-vs-All strategy that eliminates the need for a threshold to distinguish common from unknown samples [27]. In addition to the challenge of defining a threshold, CMU [24] states that relying on a single discriminability metric for OOD detection (such as the entropy used in UAN) is not sufficiently accurate and proposes a combination of metrics (entropy, consistency, and confidence). The lack of discriminability of a single metric often arises from the overconfidence exhibited by deep learning-based models [28], which limits their ability to generate uncertainty spaces for unknown samples
@@
Finally, the previously described UniDA methods were originally designed for computer vision tasks. To the best of our knowledge, the only UniDA approach tailored for TS is RAINCOAT [11], which introduces a modification of the FNO layer. In this method, an OT alignment is performed based on the variation of the target samples’ latent representations in a deep reconstruction scheme. This suggests that the presence of private class samples will modify the representation space after some optimization under a reconstruction loss. Such variation is then measured by a statistical test for bimodality on the target batch samples. The underlying RAINCOAT assumption is that: pseudo-labelling common and unknown classes is possible at the batch level. This approach is computationally expensive due to its reliance on this internal deep learning optimization step. However, some elements can be retained to approach UniDA on TS: the FNO architecture and its pseudo-labelling assumption. Using FNO allows us to build a UniDA approach tailored for TS, as in [11].
@@
In this paper, we introduce UniJDOT, a novel OT-based UniDA method for times series classification. This method can be viewed as an extension of DeepJDOT [21] to UniDA, with a focus on TS. The main novelties associated with our proposed method are:
@@
  • A joint decision space that mitigates classifier outputs using distance-based probability vectors over the feature space. This is designed to cope with the lack of discriminativity described by CMU [24].
@@
An overview of the proposed method is shown in Figure 1, and can be summarized as follows. Along with convolutional neural networks (CNNs), a layer inspired by the FNO is used to capture time-frequency features [11]. A pseudo-labelling block separates unknown and common target samples. Finally, common target classes are aligned with their source counterparts with OT, while unknown target samples are effectively isolated in a decision space. The experiments highlight the significance of each proposed module and demonstrate the superior performance of UniJDOT compared to state-of-the-art methods applied to TS.
The remainder of this paper is organized as follows: Section 2 formalizes the discrete OT, FNO-based architecture, and domain adaptation with DeepJDOT. Section 3 describes UniJDOT. Section 4 presents experimental evaluations on TS benchmarks. Finally, Section 5 concludes the paper. Our code is available at https://github.com/RomainMsrd/UniJDOT.
@@ 2.1 Fourier Neural Operator
The Fourier Neural Operator (FNO) was introduced in [10] with a focus on approximating the solution operators of partial differential equations. The FNO relies on multiple Fourier layers, which consist of projecting the output of a neural network into the Fourier space by applying a Fast Fourier Transform. Following this step, a linear transformation is applied to the lower Fourier modes while the higher modes are filtered. Finally, an Inverse Fast Fourier Transform is applied to project back the data into the original space. The original Fourier layer is slightly modified in [11], by using a cosine smoothing function to prevent alignment over noisy frequency features and extracting polar coordinates of the frequency coefficients instead of performing an inverse Fourier transform. In an abuse of notation, this modified Fourier layer will be referred to as FNO in this paper.
@@ 2.2 Discrete Optimal Transport
where $\Pi(\mathbf{a},\mathbf{b})=\{\gamma\in(\mathbb{R}^{+})^{n_{s}\times n_{t}}|\gamma\mathbbm{1}_{n_{t}}=\mathbf{a},\gamma^{\top}\mathbbm{1}_{n_{s}}=\mathbf{b}\}$ and $\mathbf{C}\in\mathbb{R}^{n_{s}\times n_{t}}$. By adding an entropic regularization to (1), this problem becomes tractable on large amounts of data using the Sinkhorn algorithm [29], thereby increasing the use of OT in data science [30].
Unbalanced Optimal Transport (UOT) [31, 32] proposes a relaxation of the mass preservation enforced by the set $\Pi(\mathbf{a},\mathbf{b})$:
@@
where $\tau_{1}$ and $\tau_{2}$ are penalization coefficients and KL is the Kullback-Leibler divergence. This formulation is deemed more robust to distribution shift and unbalanced mini-batch [22]. In our case, the masses $\mathbf{a}$ and $\mathbf{b}$ are each uniform, since there is no reason to assign different weights to individual samples. Without loss of generality, we will denote $\textbf{UOT}(\mathbf{a},\mathbf{b},\mathbf{C})$ by $\mathcal{L}_{\text{UOT}}(\mathbf{C})$ and $\textbf{OT}(\mathbf{a},\mathbf{b},\mathbf{C})$ by $\mathcal{L}_{\text{OT}}(\mathbf{C})$.
@@ 2.3 Domain Adaptation with DeepJDOT
Let $g$ be a feature extractor that maps the input space into a feature space, and $f$ be a classifier that maps the feature space to the label space. Let $h(x)=f(g(x))$. DeepJDOT [21] proposes to minimize the joint discrepancy between the distributions of each domain seen in both the feature and label spaces while optimizing the classification accuracy on the target domain. This can be done by defining the following transportation cost for each source sample $x_{i}^{s}$ and target sample $x_{j}^{t}$:
@@ 2.4 Universal Domain Adaptation
UniDA extends the scope of UDA by addressing a more general and challenging setting [12]. In UniDA, we are given a labeled source domain, $\mathcal{D}^{s}=\{(x_{i}^{s},y_{i}^{s})\}_{i=1}^{n_{s}},$ which follows a joint distribution $\mathcal{P}^{s}(x^{s},y^{s})$, and an unlabeled target domain, $\mathcal{D}^{t}=\{(x_{i}^{t})\}_{i=1}^{n_{t}},$ which follows another joint distribution $\mathcal{P}^{t}(x^{t},y^{t})$. As in UDA, it is assumed that the source and target distributions are not identical, i.e., $\mathcal{P}^{s}\neq\mathcal{P}^{t}$. UniDA operates under the flexible assumption that the label sets may differ, i.e., $\mathcal{Y}^{s}\neq\mathcal{Y}^{t}$. To capture this distinction, the label sets are divided into three subsets [12]:
@@ 3.1 Pseudo-labelling
Given a pretrained network $h$ on the source domain, our goal is to discriminate between unknown and common target data based on the classifier’s outputs. However, the classifier often exhibits overconfidence when dealing with unknown samples [28]. To mitigate this, we propose a merged approach that regularizes the classifier’s predictions by integrating distance-based probabilities within the feature space, while leveraging all target samples in a given batch. We assume that at the batch level, samples of both common and unknown classes exist. By doing so, we ensure that a relatively large distance in the feature space is associated with low model confidence.
@@
For a given batch, the maximum value of each probability vector $p_{t}^{\prime}$ is collected into an empirical scalar distribution. Such a distribution represents the prediction confidence of the model. Given our previous assumption, one can find a threshold that maximizes the separation between samples with low and high prediction confidence. Such a problem is well known in image processing, and several approaches have been developed under the so-called binary auto-thresholding (e.g. Otsu [33], Li [34, 35], Yen [36], and triangle [37]).

### v3

@@ 1 Introduction
Deep learning models have enabled significant improvements in Time Series (TS) classification, due to their powerful representational capabilities [TSsurvey]. However, most models suffer from a lack of generalization and can hardly be transferred from their training domain (the so-called source domain) to an unfamiliar, external domain (the target domain) [ADATIME]. This limitation is mainly caused by distribution shifts, a very common phenomenon in TS [ZHANG202345, GUO2023104998]. To address this problem, Unsupervised Domain Adaptation (UDA) attempts to find a domain-invariant feature space for labeled source samples and unlabeled target samples. Recently, several UDA algorithms have been benchmarked for TS tasks [ADATIME], mainly using image processing architectures. The limited performance of UDA methods on TS datasets compared to their high accuracy on image tasks underscores the need for TS-specific architectures. Since [ADATIME], many approaches have enhanced the performance of TS representation by identifying invariant temporal features specifically tailored for TS [liu2024boosting, ozyurt2023contrastive, lee2024soft, biswas2024, match_and_deform]. For example, frequency-based features extracted from neural operators, such as Fourier Neural Operators (FNO) [fno_li2020fourier], have improved domain-invariant representations and have been used for UDA on TS [RAINCOAT].
@@
The UDA framework actually covers a special case where the classes are the same in both domains (despite some small drift between domains) known as the closed-set assumption. Motivated by this limitation, Universal Domain Adaptation (UniDA) was introduced [UDA]. In UniDA, the target domain remains unlabeled, certain classes are shared between the source and target domains, while others are exclusive to either domain. Unlike standard UDA, UniDA helps to discover unknown class samples that are unique to the target domain. This capability is crucial in TS classification, where identifying unseen patterns is often critical [TSsurvey2].
@@
UniDA encompasses two primary tasks: alignment of common classes across domains and isolation of unknown Out-Of-Distribution (OOD) samples. Several approaches have been developed to address both tasks using different strategies. The alignment module is usually a pre-existing UDA method that is slightly modified to account for unknown labels. For instance, UAN [UDA] uses an adversarial UDA approach introduced in [ganin2016, tzeng2017], while DANCE [DANCE] adopts a similarity-based clustering approach inspired by [haeusser2017]. Similarly, PPOT [PPOT] and UniOT [UniOT] align source and target common samples using Optimal Transport (OT) [villani2009optimal], as proposed in [deepJDOT, JUMBOT]. Alternatively, UniAM [UniAM] introduces a novel alignment framework leveraging sparse data representation and Vision Transformer (ViT) architectures.
@@
On the other hand, the cited approaches rely on OOD detection methods that are highly threshold-dependent. These methods aim to separate common from unknown samples by thresholding discriminability metrics such as entropy [DANCE, UDA], OT masses [PPOT, UniOT], reconstruction error [UniAM], or combinations of metrics [CMU]. The threshold value is typically either a fine-tuned hyperparameter [UDA, DANCE, PPOT, UniAM] or a fixed value [UniOT] determined from a dataset. However, as shown in experiments in Section LABEL:IV, the optimal threshold often varies between datasets and across tasks. Nevertheless, most previous approaches rely on robust state-of-the-art alignment modules to address the inefficiencies in OOD detection caused by rigid threshold definitions. These approaches fail to consider that inadequate OOD detection can compromise the performance of the alignment module. In consequence, these observations highlight the need for a dynamic threshold selection to better adapt UniDA methods to different datasets.
@@
Few approaches attempt to address the dynamic nature of the threshold value. For example, TNT [TNT] computes a task-specific threshold, but this computation is based solely on the source dataset. Domain shifts may render such computations ineffective on the target dataset. Moreover, OVANet [OVANET] employs a One-vs-All strategy that eliminates the need for a threshold to distinguish common from unknown samples [padhy2020revisiting]. In addition to the challenge of defining a threshold, CMU [CMU] states that relying on a single discriminability metric for OOD detection (such as the entropy used in UAN) is not sufficiently accurate and proposes a combination of metrics (entropy, consistency, and confidence). The lack of discriminability of a single metric often arises from the overconfidence exhibited by deep learning-based models [lakshminarayanan2017simple], which limits their ability to generate uncertainty spaces for unknown samples
@@
Finally, the previously described UniDA methods were originally designed for computer vision tasks. To the best of our knowledge, the only UniDA approach tailored for TS is RAINCOAT [RAINCOAT], which introduces a modification of the FNO layer. In this method, an OT alignment is performed based on the variation of the target samples’ latent representations in a deep reconstruction scheme. This suggests that the presence of private class samples will modify the representation space after some optimization under a reconstruction loss. Such variation is then measured by a statistical test for bimodality on the target batch samples. The underlying RAINCOAT assumption is that: pseudo-labelling common and unknown classes is possible at the batch level. This approach is computationally expensive due to its reliance on this internal deep learning optimization step. However, some elements can be retained to approach UniDA on TS: the FNO architecture and its pseudo-labelling assumption. Using FNO allows us to build a UniDA approach tailored for TS, as in [RAINCOAT].
@@
In this paper, we introduce UniJDOT, a novel OT-based UniDA method for times series classification. This method can be viewed as an extension of DeepJDOT [deepJDOT] to UniDA, with a focus on TS. The main novelties associated with our proposed method are:
@@
  • A joint decision space that mitigates classifier outputs using distance-based probability vectors over the feature space. This is designed to cope with the lack of discriminativity described by CMU [CMU].
@@
An overview of the proposed method is shown in Figure 1, and can be summarized as follows. Along with convolutional neural networks (CNNs), a layer inspired by the FNO is used to capture time-frequency features [RAINCOAT]. A pseudo-labelling block separates unknown and common target samples. Finally, common target classes are aligned with their source counterparts with OT, while unknown target samples are effectively isolated in a decision space. The experiments highlight the significance of each proposed module and demonstrate the superior performance of UniJDOT compared to state-of-the-art methods applied to TS.
The remainder of this paper is organized as follows: Section 2 formalizes the discrete OT, FNO-based architecture, and domain adaptation with DeepJDOT. Section 3 describes UniJDOT. Section LABEL:IV presents experimental evaluations on TS benchmarks. Finally, Section LABEL:V concludes the paper. Our code is available at https://github.com/RomainMsrd/UniJDOT.
@@ 2.1 Fourier Neural Operator
The Fourier Neural Operator (FNO) was introduced in [fno_li2020fourier] with a focus on approximating the solution operators of partial differential equations. The FNO relies on multiple Fourier layers, which consist of projecting the output of a neural network into the Fourier space by applying a Fast Fourier Transform. Following this step, a linear transformation is applied to the lower Fourier modes while the higher modes are filtered. Finally, an Inverse Fast Fourier Transform is applied to project back the data into the original space. The original Fourier layer is slightly modified in [RAINCOAT], by using a cosine smoothing function to prevent alignment over noisy frequency features and extracting polar coordinates of the frequency coefficients instead of performing an inverse Fourier transform. In an abuse of notation, this modified Fourier layer will be referred to as FNO in this paper.
@@ 2.2 Discrete Optimal Transport
where $\Pi(\mathbf{a},\mathbf{b})=\{\gamma\in(\mathbb{R}^{+})^{n_{s}\times n_{t}}|\gamma\mathbbm{1}_{n_{t}}=\mathbf{a},\gamma^{\top}\mathbbm{1}_{n_{s}}=\mathbf{b}\}$ and $\mathbf{C}\in\mathbb{R}^{n_{s}\times n_{t}}$. By adding an entropic regularization to (1), this problem becomes tractable on large amounts of data using the Sinkhorn algorithm [sinkhorn], thereby increasing the use of OT in data science [OT_book].
Unbalanced Optimal Transport (UOT) [UOT_gilles, UOT_chizat] proposes a relaxation of the mass preservation enforced by the set $\Pi(\mathbf{a},\mathbf{b})$:
@@
where $\tau_{1}$ and $\tau_{2}$ are penalization coefficients and KL is the Kullback-Leibler divergence. This formulation is deemed more robust to distribution shift and unbalanced mini-batch [JUMBOT]. In our case, the masses $\mathbf{a}$ and $\mathbf{b}$ are each uniform, since there is no reason to assign different weights to individual samples. Without loss of generality, we will denote $\textbf{UOT}(\mathbf{a},\mathbf{b},\mathbf{C})$ by $\mathcal{L}_{\text{UOT}}(\mathbf{C})$ and $\textbf{OT}(\mathbf{a},\mathbf{b},\mathbf{C})$ by $\mathcal{L}_{\text{OT}}(\mathbf{C})$.
@@ 2.3 Domain Adaptation with DeepJDOT
Let $g$ be a feature extractor that maps the input space into a feature space, and $f$ be a classifier that maps the feature space to the label space. Let $h(x)=f(g(x))$. DeepJDOT [deepJDOT] proposes to minimize the joint discrepancy between the distributions of each domain seen in both the feature and label spaces while optimizing the classification accuracy on the target domain. This can be done by defining the following transportation cost for each source sample $x_{i}^{s}$ and target sample $x_{j}^{t}$:
@@ 2.4 Universal Domain Adaptation
UniDA extends the scope of UDA by addressing a more general and challenging setting [UDA]. In UniDA, we are given a labeled source domain, $\mathcal{D}^{s}=\{(x_{i}^{s},y_{i}^{s})\}_{i=1}^{n_{s}},$ which follows a joint distribution $\mathcal{P}^{s}(x^{s},y^{s})$, and an unlabeled target domain, $\mathcal{D}^{t}=\{(x_{i}^{t})\}_{i=1}^{n_{t}},$ which follows another joint distribution $\mathcal{P}^{t}(x^{t},y^{t})$. As in UDA, it is assumed that the source and target distributions are not identical, i.e., $\mathcal{P}^{s}\neq\mathcal{P}^{t}$. UniDA operates under the flexible assumption that the label sets may differ, i.e., $\mathcal{Y}^{s}\neq\mathcal{Y}^{t}$. To capture this distinction, the label sets are divided into three subsets [UDA]:
@@ 3.1 Pseudo-labelling
Given a pretrained network $h$ on the source domain, our goal is to discriminate between unknown and common target data based on the classifier’s outputs. However, the classifier often exhibits overconfidence when dealing with unknown samples [lakshminarayanan2017simple]. To mitigate this, we propose a merged approach that regularizes the classifier’s predictions by integrating distance-based probabilities within the feature space, while leveraging all target samples in a given batch. We assume that at the batch level, samples of both common and unknown classes exist. By doing so, we ensure that a relatively large distance in the feature space is associated with low model confidence.
@@
For a given batch, the maximum value of each probability vector $p_{t}^{\prime}$ is collected into an empirical scalar distribution. Such a distribution represents the prediction confidence of the model. Given our previous assumption, one can find a threshold that maximizes the separation between samples with low and high prediction confidence. Such a problem is well known in image processing, and several approaches have been developed under the so-called binary auto-thresholding (e.g. Otsu [otsu], Li [li_thr, li_thr_iterative], Yen [yen_thresh], and triangle [triangle_method]).